OpenStudy (anonymous):

Show all work in simplifying the quantity 3x^2 - 3x - 60 / x^2 +9x + 20 (Be sure to list restrictions).

OpenStudy (callisto):

3x^2 - 3x - 60 = (x-5) (x+4) x^2 +9x + 20 = (x+4)(x+5) 3x^2 - 3x - 60 / x^2 +9x + 20 = (x-5) (x+4) / (x+4)(x+5) = (x-5) /(x+5) for it to exist x=/= -5 and x=/=-4

OpenStudy (anonymous):

what is =/= ?

OpenStudy (anonymous):

\[(3x^2 - 3x - 60) / (x^2 +9x + 20 )=3(x^2-x-20)/(x+4)(x+5)\]\[=3(x+4)(x-5)/(x+4)(x+5)=3(x-5)/(x+5)\]\[x \neq-4,x \neq-5\]

OpenStudy (anonymous):

Callisto, I think you dropped a factor of three in the numerator.

OpenStudy (callisto):

yup , sorry!